User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Ho Module: JIVE-064, start pos 0 GB 4:44UT - small power glitch caused Mark 4 rack to reset some values, setupsx was run again and fixed the problem (Imogen)

Some how the log was set to station.log at 27.13:19, and most of the experiment was recorded to that log file. Copied out the appropriate for final log.

Hb Module: BKG+0133, start pos 0 GB

Ke Module:USN-0131, start pos 0 GB

DBBC not behaving, so couldn't observe (Ellen)

Yg Module: USN-0175, start pos 0 GB

Jamie tried to fix the power levels but couldn't. As a results, tsys bands 05 -12 are all bad. (Ellen)

You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/crds75.txt · Last modified: 2015/01/28 18:00 by David Horsley